OpenStudy (anonymous):

Quick exponent rule question Everyone knows that if you are deviding two terms that have the same bases you subtract the exponents. Like Z^3/Z^1 = Z^2 However well covering a basic prime factoring subject in a math book the book gives the exaple of an expression being prime factored with the variable Z in the numerator and Z^3 in the denominator and calls it Z^2 How is that possible? Why is it not Z^-2 The z ontop is obviously carries the expontent 1 and last i check 1 - 3 = -2 Someone please explain

OpenStudy (anonymous):

Z / Z^3 = 1 / Z^2 or Z^-2

OpenStudy (anonymous):

care to elaborate

OpenStudy (anonymous):

how does z/Z^3 = 1?

OpenStudy (anonymous):

it doesnt. he said it = 1/z^2 or the negative recip of z^-2

OpenStudy (anonymous):

ok so intead of righting the z^-2 the choose to right it as a positive exponent as 1/z^2

OpenStudy (anonymous):

makes sense i guess, i wish the book would explain that instead of assuming you know what its doing